OpenStudy (anonymous):

The equation of line EF is y = x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, -2). A.y = -2x - 2 B. y = x - 2 C. y = x - 2 D. y = 2x - 2

OpenStudy (solomonzelman):

Ok, so what's going to change/differ in the line parallel to EF, is the y-intercept, but the slope will remain the same.

OpenStudy (solomonzelman):

y=x+b we know that slope is still one, but what is the intercept? Plug in your point, (0,-2) to find b. -2=(0)+b -2=b So, y=x-2 is the equation of a line parallel to EF, that passes through the point (0,-2)